# Towards Testing and Evaluating Vision-Language-Action Models for Robotic Manipulation: An Empirical Study

### v1

@@ 7. Related Work
Foundation models for robotics. A large body of research has been done on the use of foundation models for robotics. Based on the data modality, recent work can be roughly categorized into two groups: (1) those that use large language models (LLMs) for robotics, and (2) those that use multi-modality foundation models (e.g., vision language models) for robotics. One of the early attempts in using LLMs for robotics is to leverage LLMs as a reward designer for Deep Reinforcement Learning-based robotic manipulation (Song et al., 2023b). This work proposes a self-refined framework to let the LLM automatically and interactively generate and refine reward functions used to train DRL policies for robotic control. Alternatively, Zhou et al. (Zhou et al., 2024) adopts LLMs, first, as a translator to convert natural language input into a Planning Domain Definition Language (PDDL) formulation and then outputs action plan sequences for hong-horizon tasks.
@@
Different from the use of LLMs, multi-modality foundation models such as vision language models (VLMs) enable the possibility of advancing AI-enabled robotics in several novel tasks, e.g., for robotic manipulation (Brohan et al., 2022, 2023; Team et al., 2024), visual question answering with robots (Du et al., 2023), and visual state representations (Nair et al., 2022). Our work is most closely related to those using vision-language-action models, a specialized category of VLMs, for robotic manipulation (Brohan et al., 2023; Team et al., 2024; Kim et al., 2024; Li et al., 2024c; Brohan et al., 2022; Li et al., 2023; Stone et al., 2023). One of the pioneer works in VLA models for robotic manipulation is RT-1 (Brohan et al., 2022), which uses a combination of a FiLM EfficientNet and a transformer to learn control policies from 130k real-world robot demonstrations. Since the publish of Open X-Embodiment dataset (Padalkar et al., 2023), a series of VLA models have been proposed by either training or fine-tuning on Open X-Embodiment (Brohan et al., 2023; Team et al., 2024; Kim et al., 2024; Li et al., 2024c). Our work is parallel to these works. We propose a general testing framework to test VLA models and present an empirical study to comprehensively assess the performance of VLA models.
@@
Testing cyber-physical systems. Quality assurance of traditional cyber-physical systems (CPS) and AI-driven CPS is a vital topic in the realm of software engineering, which assures the safety and trustworthiness of such complex systems in safety-critical domains. Substantial efforts have been made by researchers and industrial practitioners to safeguard the quality of CPS from various perspectives, such as testing (Lee et al., 2023; Menghi et al., 2020; Chen et al., 2020; Ayerdi et al., 2021; Zhang et al., 2022), analysis (Stocco and Tonella, 2022; Biagiola and Tonella, 2024; Yohanandhan et al., 2020) and repairing (Song et al., 2023a; Wang et al., 2024; Guo et al., 2020). Considering robotic manipulation as a widely deployed representative CPS, our work is most related to the testing of CPS.
@@
In particular, Lee et al. (Lee et al., 2023) propose MOTIF, a gray-box fuzzing-based approach to generate test data for software deployed in CPS. This method monitors the coverage achieved by the vanilla and the mutated functions, thereby exploring the behaviour space of the software under test. Differing from the existing testing solutions, MOTIF specifically targets the testing challenges of C and C++ languages that are widely used in CPS contexts. In terms of the black-box approach, Menghi et al. (Menghi et al., 2020) introduces ARIsTEO, a searched-based testing method with a loop of approximation and refinement to identify requirement violations of CPS models, and Chen et al. (Chen et al., 2020) propose an active fuzzing approach to find test suits for packet-level CPS network attacks. Moving forward with CPS with AI module embedded, Zhang et al. (Zhang et al., 2022) design a series of time-aware coverage criteria for DNN controllers in CPS and develop a falsification framework FalsifAI, to find system defects by leveraging the coverage information from the proposed criteria. It is worth noting that the aforementioned works are not well applicable to VLA models, which is attributable to the multimodality characteristic, autoregressive generation mechanism, and large-scale model size. In contrast, our study initiates a testing framework tailored for VLA models in the context of different robotic manipulation tasks. VLATest encompasses four multimodality-aware testing operators to assess the VLA models’ performance from different perspectives.
@@
Benchmarking and testing foundation models. After the remarkable success of foundation models, an important research direction is to explore and understand the boundaries of their capabilities, which can provide a basis for further safeguarding and enhancement. Based on this motivation, numerous studies have focused on evaluating large language models in the text domain across various properties, such as correctness (Liang et al., 2023a), factuality (Lin et al., 2022; Dziri et al., 2022), robustness (Zhu et al., 2023; Wang et al., 2023), fairness (Sun et al., 2024), and privacy (Hamid et al., 2023). These studies offer a wealth of resources for the general performance assessment of LLMs. On the other hand, researchers have also made significant efforts to evaluate specific capabilities of LLMs, such as code-related abilities. These efforts include validating the correctness of generated code (Cassano et al., 2023; Liu et al., 2024b; Yu et al., 2024), examining whether LLMs can address real-world GitHub issues (Jimenez et al., 2023), and analyzing error patterns in LLM-driven code translations (Pan et al., 2024).
@@ 8. Threats to Validity
Internal Validity. One potential threat comes from the randomness in our experiments. To mitigate this, we generated 18,604 testing scenes in our empirical study. Overall, our empirical study took over 400 GPU hours to finish. While we only executed each testing scene once per VLA model, we believe this has limited impact on our findings as our goal is to evaluate the VLA model’s overall performance on each task instead of the success of one specific testing scene.
Another potential threat to internal validity is the choice of prompt templates. To mitigate this, we followed the previous works (Team et al., 2024; Kim et al., 2024; Brohan et al., 2022, 2023) to use the standard prompt template when evaluating VLA models for robotic manipulation. We also discuss the impact of prompts and the potential of improving VLA models with better prompting in Sec. 6.
External Validity. Threats to external validity include whether our study results can generalize to different experimental settings, such as different robotic manipulation tasks and different VLA models. To mitigate these threats, we select the four most popular robotic manipulation tasks according to Open X-Embodiment dataset (Padalkar et al., 2023). In terms of the VLA models, we included the SOTA publicly available VLA models by the time we conducted this study. Given that the VLA model for robotic manipulation is a fast-developing research area, we also plan to extend our evaluation to other SOTA VLA models, e.g., RT-2 (Brohan et al., 2023), once they are made available.
Finally, we have only implemented VLATest and evaluated the VLA models within one simulation environment, Maniskill2 (Gu et al., 2023). To reduce the impact of distribution shifts between simulation and real robot execution, we chose a improved version of Maniskill2 (Li et al., 2024b), which provides visual matching to render realistic images as the visual input for VLA models.
Construct Validity. When designing our testing framework, VLATest, we have only considered a limited number of testing operators. To combat the threat, we included testing operators covering different aspects, e.g., task difficulties (i.e., number of obstacles) and environmental factors (i.e., the lighting intensity and camera poses). In future work, one may continue to improve VLATest by adding more testing operators, such as the number of lighting sources, the intrinsic parameters of cameras, and the resolution of cameras, for more comprehensive evaluation of VLA models.
@@ 9. Conclusion
In this paper, we propose VLATest, one of the early testing frameworks for testing VLA models for robotic manipulation. VLATest is a generation-based fuzzing framework based on ten operators covering different perspectives in the test scene of robotic manipulation. Upon implementing VLATest with Maniskill2 simulation environments, we further conducted a large-scale empirical study to assess the performance and robustness of seven popular VLA models across four robotic manipulation tasks. We generated 18,604 testing scenes, conducted more than 400 GPU hours of simulation and performed a detailed analysis of the challenges and limitations faced by the current VLA models. At the end of the paper, we discuss the implications of our study, shedding light on several future research directions to improve the quality and reliability of VLA models for robotic manipulation.

### v2

@@ 7. Related Work
Foundation Models for Robotics. A large body of research has been done on the use of foundation models for robotics. Based on the data modality, recent work can be roughly categorized into two groups: (1) those that use LLMs for robotics and (2) those that use multi-modality foundation models (e.g., VLMs) for robotics. One of the early attempts in using LLMs for robotics is to leverage LLMs as a reward designer for Deep Reinforcement Learning-based robotic manipulation (Song et al., 2023b). This work proposes a self-refined framework to let the LLM automatically generate and refine reward functions used to train policies for robotic control. Alternatively, Zhou et al. (Zhou et al., 2024) adopts LLMs to convert natural language input into a Planning Domain Definition Language formulation before outputting action plan sequences.
@@
Different from the use of LLMs, multi-modality foundation models such as VLMs enable the possibility of advancing AI-enabled robotics in several novel tasks, e.g., for robotic manipulation (Brohan et al., 2022, 2023; Team et al., 2024), visual question answering with robots (Du et al., 2023), and visual state representations (Nair et al., 2023). Our work is most closely related to those using vision-language-action models, a specialized category of VLMs, for robotic manipulation (Brohan et al., 2023; Team et al., 2024; Kim et al., 2024; Li et al., 2025; Brohan et al., 2022; Li et al., 2024c; Stone et al., 2023). One of the pioneer works in VLA models for robotic manipulation is RT-1 (Brohan et al., 2022), which uses a combination of a FiLM EfficientNet and a transformer to learn control policies from 130k real-world robot demonstrations. Since the release of the Open X-Embodiment dataset (Padalkar et al., 2023), a series of VLA models have been proposed by either training or fine-tuning on Open X-Embodiment (Brohan et al., 2023; Team et al., 2024; Kim et al., 2024; Li et al., 2025). Our work is parallel to these works. We propose a general testing framework to test VLA models and present an empirical study to comprehensively assess the performance of VLA models.
@@
Testing Cyber-Physical Systems. Quality assurance of traditional CPS and AI-driven CPS is a vital topic in software engineering, as it ensures the safety and trustworthiness of such complex systems in safety-critical domains. Substantial efforts have been made by researchers and industrial practitioners to safeguard the quality of CPS from various perspectives, such as testing (Lee et al., 2023; Menghi et al., 2020; Chen et al., 2020; Ayerdi et al., 2021; Zhang et al., 2022), analysis (Stocco and Tonella, 2022; Biagiola and Tonella, 2024; Yohanandhan et al., 2020) and repairing (Song et al., 2023a; Wang et al., 2024; Guo et al., 2020). Considering robotic manipulation as a widely deployed representative CPS, our work is most related to the testing of CPS.
@@
In particular, Lee et al. (Lee et al., 2023) propose MOTIF, a gray-box fuzzing-based approach to generate test data for software deployed in CPS. This method monitors the coverage achieved by the vanilla and the mutated functions, thereby exploring the behavior space of the software under test. Differing from the existing testing solutions, MOTIF specifically targets the testing challenges of C and C++ languages that are widely used in CPS contexts. In terms of the black-box approach, Menghi et al. (Menghi et al., 2020) introduce ARIsTEO, a search-based testing method with a loop of approximation and refinement to identify requirement violations of CPS models, and Chen et al. (Chen et al., 2020) propose an active fuzzing approach to find test suites for packet-level CPS network attacks. Moving forward with CPS with an AI module embedded, Zhang et al. (Zhang et al., 2022) design a series of time-aware coverage criteria for DNN controllers in CPS and develop a falsification framework, FalsifAI, to find system defects by leveraging the coverage information from the proposed criteria. It is worth noting that the aforementioned works are not well applicable to VLA models, which is attributable to the multimodality characteristic, the autoregressive generation mechanism, and the large-scale model size. In contrast, our study initiates a testing framework tailored for VLA models in the context of different robotic manipulation tasks. VLATest encompasses four multimodality-aware testing operators to comprehensively assess the VLA models from different perspectives.
@@
Benchmarking and Testing Foundation Models. After the remarkable success of foundation models, an important research direction is to explore and understand the boundaries of their capabilities, which can provide a basis for further safeguarding and enhancement. Based on this motivation, numerous studies have focused on evaluating large language models in the text domain across various properties, such as correctness (Liang et al., 2023a), factuality (Lin et al., 2022; Dziri et al., 2022), robustness (Zhu et al., 2024; Wang et al., 2023), fairness (Sun et al., 2024), and privacy (Hamid et al., 2023). These studies offer a wealth of resources for the general performance assessment of LLMs. On the other hand, researchers have also made significant efforts to evaluate specific capabilities of LLMs, such as code-related abilities. These efforts include validating the correctness of generated code (Cassano et al., 2023; Liu et al., 2024b; Yu et al., 2024), examining whether LLMs can address real-world GitHub issues (Jimenez et al., 2024), and analyzing error patterns of LLM-generated code (Pan et al., 2024; Wang et al., 2025).
@@ 8. Threats to Validity
Internal Validity. One potential threat comes from the randomness of our experiments. To mitigate this, we generated 18,604 testing scenes in our empirical study. Overall, our empirical study took over 580 GPU hours to finish. While we only executed each testing scene once per VLA model, we believe this has limited impact on our findings as our goal is to evaluate the VLA model’s overall performance on each task instead of the success of one specific testing scene.
Another potential threat to internal validity is the choice of prompt templates. To mitigate this, we followed the previous works (Team et al., 2024; Kim et al., 2024; Brohan et al., 2022, 2023) to use the standard prompt template when evaluating VLA models for robotic manipulation in RQ1 ~ RQ5. In RQ6, we evaluated the VLA models’ performance with the mutated prompt templates. However, the prompt templates mutated by GPT-4o may not carry the same meanings as the original ones. To mitigate the threat, two of the authors manually validated that the mutated prompts were semantically equivalent to the corresponding standard prompts. In addition, we discuss the impact of prompts and the potential of improving VLA models with better prompting in Sec. 6.
External Validity. Threats to external validity include whether our study results can generalize to different experimental settings, such as different robotic manipulation tasks and different VLA models. To mitigate these threats, we select the four most popular robotic manipulation tasks according to the Open X-Embodiment dataset (Padalkar et al., 2023). In terms of the VLA models, we included the SOTA publicly available VLA models by the time we conducted this study. Given that the VLA model for robotic manipulation is a fast-developing research area, we also plan to extend our evaluation to other SOTA VLA models, e.g., RT-2 (Brohan et al., 2023), once they are made available.
Finally, we have only implemented VLATest and evaluated the VLA models within one simulation environment, Maniskill2 (Gu et al., 2023). To reduce the impact of distribution shifts between simulation and real robot execution, we chose an improved version of Maniskill2 (Li et al., 2024b), which provides visual matching to render realistic images as the visual input for VLA models.
Construct Validity. When designing our testing framework, VLATest, we have only considered a limited number of testing operators. To combat the threat, we included testing operators covering different aspects, e.g., task difficulties (i.e., number of confounding objects) and environmental factors (i.e., the lighting intensity and camera poses). In future work, one may continue to improve VLATest by adding more testing operators, such as the number of lighting sources, the intrinsic parameters of cameras, and the resolution of cameras, for a more comprehensive evaluation of VLA models. Another threat to construct validity is the coverage metric we used in RQ1. Although our trajectory coverage metric may not be the best metric for measuring test sufficiency, we believe it is a pragmatic choice since there is no well-established coverage metric for either VLA models or robotic manipulation. In future work, one may consider designing a new coverage metric given the special characteristics of robotic manipulation with VLA models.
@@ 9. Conclusion
In this paper, we propose VLATest, one of the early testing frameworks for testing VLA models for robotic manipulation. VLATest is a generation-based fuzzing framework based on ten operators covering different perspectives in the test scene of robotic manipulation. Upon implementing VLATest with Maniskill2 simulation environments, we further conducted a large-scale empirical study to assess the performance and robustness of seven popular VLA models across four robotic manipulation tasks. We generated 18,604 testing scenes, conducted more than 580 GPU hours of simulation, and performed a detailed analysis of the challenges and limitations faced by the current VLA models. At the end of the paper, we discuss the implications of our study, shedding light on several future research directions to improve the quality and reliability of VLA models.
10. Data Availability
Our artifacts, including the replication packages and the generated testing scenes, are available on a Git repository: https://github.com/ma-labo/VLATest.
Acknowledgements.
This work was supported in part by the Canada CIFAR AI Chairs Program, Natural Sciences and Engineering Research Council of Canada, JST CRONOS Grant (No. JPMJCS24K8), JSPS KAKENHI Grant (No. JP21H04877, No. JP23H03372, and No. JP24K02920), and the Autoware Foundation.